# GI-NAS: Boosting Gradient Inversion Attacks through Adaptive Neural Architecture Search

### v2

@@ IV-A2 Optimal Model Selection
where $\mathcal{L}_{grad}(\cdot)$ is the gradient matching loss, $\mathcal{T}(\cdot)$ is the estimated gradient transformation [16], and $\mathbf{\phi}_{r}$ are the parameters of $G_{r}$. Here we introduce $\mathcal{T}(\cdot)$ to estimate the gradient transformation following the previous defense auditing work [16], since the victims may apply defense strategies to $\mathbf{g}$ (e.g., Gradient Clipping [30]) and only release disrupted forms of gradients. Empirically, the model that can perform the best at the start and stand out from numerous candidates is likely to have better implicit architectural priors with respect to the private batch. Therefore, we regard the model that achieves the minimal initial loss as our best model $G_{opt}$ and update our selection during the traversal. Since only the initial loss is calculated and no back-propagation is involved, this search process is training-free and hence computationally efficient.
@@ V-A Experimental Setup
State-of-the-art Baselines for Comparison. We implement the following gradient inversion baselines: (1) IG (Inverting Gradients) [14]: pixel-level reconstruction with angle-based loss function; (2) GI (GradInversion) [15]: realizing batch-level restoration via multiple regularization priors; (3) GGL (Generative Gradient Leakage) [16]: employing strong GAN priors to produce high-fidelity images under severe defense strategies; (4) GIAS (Gradient Inversion in Alternative Space) [17]: searching the optimal latent code while optimizing in the generator parameter space; (5) GION (Gradient Inversion via Over-parameterized Networks) [21]: designing an over-parameterized convolutional network with excessive parameters and employing a fixed network architecture as implicit regularization. Note that when implementing GAN-based methods such as GGL and GIAS, we adopt BigGAN [77] pre-trained on ImageNet, which may result in mismatched priors when the target data is from CIFAR-10. However, such mismatch can be very common in realistic FL scenarios.
Quantitative Metrics. We utilize four metrics to measure the quality of reconstruction images: (1) PSNR $\uparrow$ (Peak Signal-to-Noise Ratio); (2) SSIM $\uparrow$ (Structural Similarity Index Measure); (3) FSIM $\uparrow$ (Feature Similarity Index Measure); (4) LPIPS $\downarrow$ (Learned Perceptual Image Patch Similarity) [78]. Note that $\downarrow$” indicates that the lower the metric, the better the attack performance while $\uparrow$” indicates that the higher the metric, the better the attack performance.
@@ V-B Choices of search size $n$ and network depth $t$
To obtain the best reconstruction performance, we need to carefully decide on the values of search size $n$ and network depth $t$. Note that both these two hyper-parameters reflect the degree of NAS, and there is a trade-off between under-fitting and over-fitting when choosing their values. Thus, we randomly select images (strictly having no intersections with the images used in the remaining experiments of this paper) from ImageNet and test the attack results under different values of $n$ and $t$ with the default batch size $B=4$. From Table I and Table II, we find that when $n$ and $t$ are small, increasing any one of them indeed improves the attacks, as more optimal alternatives are provided or the models are more complicated to tackle the under-fitting. However, when $n$ and $t$ are large, further increasing any one of them cannot improve the attacks and will conversely lead to performance degradation due to the over-fitting. This is because candidates that focus too much on the minimal initial loss while neglecting the fundamental patterns of data recovery may be included in the search space and eventually selected. Thus, we adopt $n=5000$ and $t=5$ in the remaining experiments of this paper, as they strike a better balance and acquire the best results in Table I and Table II.
@@
Dataset Metric IG GI GGL GIAS GION GI-NAS CIFAR-10 PSNR $\uparrow$ 16.3188 15.4613 12.4938 17.3687 30.8652 35.9883 SSIM $\uparrow$ 0.5710 0.5127 0.3256 0.6239 0.9918 0.9983 FSIM $\uparrow$ 0.7564 0.7311 0.6029 0.7800 0.9960 0.9991 LPIPS $\downarrow$ 0.4410 0.4878 0.5992 0.4056 0.0035 0.0009 ImageNet PSNR $\uparrow$ 7.9419 8.5070 11.6255 10.0602 21.9942 23.2578 SSIM $\uparrow$ 0.0815 0.1157 0.2586 0.2408 0.6188 0.6848 FSIM $\uparrow$ 0.5269 0.5299 0.5924 0.5719 0.8198 0.8513 LPIPS $\downarrow$ 0.7194 0.7168 0.6152 0.6563 0.4605 0.3952
V-C Comparison with State-of-the-art Methods
Firstly, we compare GI-NAS with state-of-the-art gradient inversion methods on CIFAR-10. From Table III, we conclude that GI-NAS achieves the best results with significant performance improvement. For instance, we realize a $5.12$ dB PSNR increase than GION, and our LPIPS value is $74.3\%$ smaller than that of GION. These prove that in contrast to GION that optimizes on a fixed network, our NAS strategy indeed comes into effect and better leverages the implicit architectural priors.
@@ V-C Comparison with State-of-the-art Methods
Dataset Metric Batch Size IG GI GGL GIAS GION GI-NAS Batch Size IG GI GGL GIAS GION GI-NAS CIFAR-10 PSNR $\uparrow$ 16 10.2647 11.2593 9.9964 11.1952 24.8746 30.5315 32 9.5802 10.5863 9.4104 9.8669 28.6832 33.0586 SSIM $\uparrow$ 0.2185 0.2431 0.1658 0.2638 0.9816 0.9948 0.1595 0.2020 0.1481 0.1724 0.9892 0.9974 FSIM $\uparrow$ 0.5254 0.5350 0.4988 0.5569 0.9869 0.9965 0.5186 0.5325 0.4995 0.5180 0.9882 0.9984 LPIPS $\downarrow$ 0.6254 0.6225 0.6655 0.5804 0.0177 0.0035 0.6598 0.6626 0.6922 0.6362 0.0218 0.0017 PSNR $\uparrow$ 48 9.5446 10.2000 9.1200 9.6376 29.5312 38.7054 96 9.2363 9.9500 8.8575 9.2982 28.9723 31.8026 SSIM $\uparrow$ 0.1574 0.1697 0.1552 0.1652 0.9932 0.9983 0.1402 0.1643 0.1516 0.2180 0.9878 0.9941 FSIM $\uparrow$ 0.5112 0.5167 0.4869 0.5126 0.9949 0.9987 0.4961 0.5238 0.4801 0.5365 0.9865 0.9941 LPIPS $\downarrow$ 0.6660 0.6571 0.7213 0.6554 0.0072 0.0020 0.6865 0.6593 0.7248 0.7297 0.0227 0.0071 ImageNet PSNR $\uparrow$ 8 7.6011 8.3419 11.5454 9.6216 20.4841 20.8451 16 7.7991 8.0233 11.4766 9.2563 20.3078 21.4267 SSIM $\uparrow$ 0.0762 0.1113 0.2571 0.2218 0.5681 0.6076 0.0704 0.0952 0.2561 0.2189 0.5507 0.6141 FSIM $\uparrow$ 0.5103 0.5365 0.5942 0.5765 0.7908 0.8105 0.5077 0.5165 0.5872 0.5832 0.7840 0.8157 LPIPS $\downarrow$ 0.7374 0.7251 0.6152 0.6621 0.5162 0.4613 0.7424 0.7362 0.6165 0.6656 0.5313 0.4613 PSNR $\uparrow$ 24 7.1352 7.9079 10.9910 9.6601 20.6845 21.7933 32 7.1085 7.8752 10.8987 10.0563 16.1275 20.0011 SSIM $\uparrow$ 0.0549 0.0892 0.2583 0.2270 0.5868 0.6236 0.0554 0.0763 0.2558 0.2361 0.3415 0.5485 FSIM $\uparrow$ 0.4794 0.4943 0.5811 0.5618 0.7978 0.8163 0.4940 0.4951 0.5889 0.5762 0.6823 0.7819 LPIPS $\downarrow$ 0.7540 0.7475 0.6261 0.6667 0.5219 0.4651 0.7493 0.7427 0.6296 0.6674 0.6291 0.5265
@@
Dataset Metric Defense IG GI GGL GIAS GION GI-NAS Defense IG GI GGL GIAS GION GI-NAS CIFAR-10 PSNR $\uparrow$ Gaussian Noise 8.8332 8.2797 10.3947 9.8146 9.7658 10.5192 Gradient Sparsification 9.6183 10.9486 10.4323 12.1328 30.3518 39.8739 SSIM $\uparrow$ 0.1177 0.0976 0.1856 0.1301 0.0280 0.0375 0.1592 0.2168 0.1862 0.3227 0.9920 0.9987 FSIM $\uparrow$ 0.4844 0.4806 0.5123 0.4843 0.5945 0.6300 0.5151 0.5329 0.5223 0.6015 0.9951 0.9987 LPIPS $\downarrow$ 0.7258 0.7077 0.6255 0.7129 0.6752 0.6740 0.6895 0.6855 0.6298 0.5582 0.0047 0.0014 PSNR $\uparrow$ Gradient Clipping 16.6203 11.5888 10.1373 17.9859 32.4744 36.2490 Soteria 9.4103 10.2939 10.4217 15.5825 31.1003 34.1804 SSIM $\uparrow$ 0.5952 0.2851 0.1760 0.6703 0.9933 0.9985 0.1404 0.1707 0.1889 0.5405 0.9900 0.9973 FSIM $\uparrow$ 0.7464 0.4976 0.5070 0.7751 0.9966 0.9992 0.5086 0.5086 0.5144 0.7133 0.9952 0.9986 LPIPS $\downarrow$ 0.3811 0.6413 0.6205 0.3054 0.0029 0.0008 0.7008 0.6940 0.6335 0.3748 0.0038 0.0014 ImageNet PSNR $\uparrow$ Gaussian Noise 7.7020 7.4553 9.7381 6.8520 9.2449 11.2567 Gradient Sparsification 8.6540 8.6519 9.8019 9.7894 13.7358 15.4331 SSIM $\uparrow$ 0.0311 0.0197 0.2319 0.1045 0.0277 0.0416 0.0669 0.0627 0.2315 0.2515 0.2387 0.3154 FSIM $\uparrow$ 0.4518 0.3804 0.5659 0.4562 0.5714 0.5846 0.5073 0.4696 0.5624 0.5664 0.6365 0.6786 LPIPS $\downarrow$ 0.7775 0.8220 0.6639 0.7282 0.8033 0.7805 0.7480 0.7681 0.6589 0.6604 0.6909 0.6535 PSNR $\uparrow$ Gradient Clipping 8.0690 7.2342 11.3215 9.7556 22.2016 23.2074 Soteria 6.5675 6.6693 11.5690 9.7588 22.9617 23.7917 SSIM $\uparrow$ 0.0881 0.0659 0.2531 0.2547 0.6266 0.7023 0.0323 0.0382 0.2623 0.2425 0.6698 0.7143 FSIM $\uparrow$ 0.5362 0.5434 0.5900 0.5794 0.8229 0.8635 0.4929 0.4454 0.5964 0.5697 0.8430 0.8658 LPIPS $\downarrow$ 0.7227 0.7230 0.6117 0.6485 0.4603 0.3743 0.7506 0.7703 0.6031 0.6631 0.4253 0.3655
@@ V-D Further Analysis
Effectiveness of Our Training-Free Search Metric. We provide extensive empirical evidence that our training-free search metric (i.e., the initial gradient matching loss) highly negatively correlates with the real attack performance. Specifically, we attack the same private batch using models with different architectures and report the PSNR performance as well as the training-free search metric of each model both on CIFAR-10 and ImageNet. Following previous training-free NAS methods [80, 81, 82], we calculate the correlation indicators (i.e., Kendall’s $\tau$ [83], Pearson Coefficient [84], Spearman Coefficient [85]) to quantitatively measure the relevance. To be more convincing, we introduce Gaussian Noise $\mathcal{N}\sim(0,1)$ and Uniform Noise $\mathcal{U}\sim(0,1)$ for comparison. From Figure 5 and Table VI, we conclude that a smaller initial gradient matching loss is more likely to result in better PSNR performance, while the random noises are far from similar effects. Thus, utilizing the initial gradient matching loss as the training-free search metric is reasonable and meaningful.
@@
Dataset Metric IG GI GGL GIAS GION GI-NAS CIFAR-10 NAS Time - - - - - 6.6 Total Time 46.1 32.9 131.4 214.8 46.9 57.3 PSNR 16.3188 15.4613 12.4938 17.3687 30.8652 35.9883 ImageNet NAS Time - - - - - 10.9 Total Time 291.1 273.7 244.6 343.2 58.1 84.4 PSNR 7.9419 8.5070 11.6255 10.0602 21.9942 23.2578
@@
TABLE VIII: Results when attacking the same batch by different networks using different parameter initialization seeds on CIFAR-10 ($32\times 32$) and ImageNet ($256\times 256$) with the default batch size $B=4$. Note that $\sigma_{loss}$ and $\sigma_{psnr}$ are respectively the standard deviations of the initial gradient matching losses and PSNR results on each network.
Dataset Model Random Seed Initial Gradient Matching Loss (×$\times 10^{-3}$) $\sigma_{loss}$ (×$\times 10^{-4}$) PSNR $\sigma_{psnr}$ CIFAR-10 $G_{c1}$ $s_{c1}$ 8.5647 0.0012 28.7283 0.1211 $s_{c2}$ 8.5648 28.8082 $s_{c3}$ 8.5645 28.5703 $G_{c2}$ $s_{c1}$ 8.4321 0.0006 33.5848 0.2581 $s_{c2}$ 8.4323 33.4631 $s_{c3}$ 8.4322 33.9584 $G_{c3}$ $s_{c1}$ 7.9501 0.0015 49.4290 0.2936 $s_{c2}$ 7.9499 49.1171 $s_{c3}$ 7.9502 49.7039 ImageNet $G_{i1}$ $s_{i1}$ 21.0167 5.6866 22.9013 0.0329 $s_{i2}$ 22.1401 22.8912 $s_{i3}$ 21.4250 22.8398 $G_{i2}$ $s_{i1}$ 24.4600 2.2089 21.1475 0.3040 $s_{i2}$ 24.0722 21.6449 $s_{i3}$ 24.0828 21.6989 $G_{i3}$ $s_{i1}$ 18.5319 6.0290 23.5278 0.1884 $s_{i2}$ 18.3819 23.2974 $s_{i3}$ 17.4208 23.6709
Robustness to Network Parameters Initialization. Following the general paradigm of previous training-free NAS methods [80, 81, 82] in other areas, the parameters of all the NAS candidates are randomly initialized and there is no special need to tune the parameters before calculating the initial gradient matching losses. Moreover, GI-NAS records both the architectures as well as the initial parameters when finding the optimal models. Therefore, for the model we decide on, its initial gradient matching loss in the NAS time (i.e., Stage 1 as shown in Figure 2) is the same as its initial gradient matching loss in the optimization time (i.e., Stage 2 as shown in Figure 2), and the influences of parameters initialization can be further reduced. To further demonstrate the robustness, we randomly select models with different architectures and apply different parameter initialization seeds to them to attack the same batch. From Table VIII, we find that for the same architecture, both the initial gradient matching losses and the final PSNR results stay within a steady range when changing the parameter initialization seeds. However, when the architectures change, both the initial gradient matching losses and the final PSNR results will vary significantly. The impacts of different architectures are much greater than the impacts of different parameter initialization seeds. As a result, we conclude that GI-NAS is stable and insensitive to network parameters initialization. We also observe that the models with smaller initial gradient matching losses often have better final PSNR results in Table VIII, which is consistent with our aforementioned findings in Figure 5 and Table VI. This again validates the rationality and usefulness of adopting the initial gradient matching loss as the training-free search metric.
Robustness to Latent Codes Initialization. The latent code $\mathbf{z_{0}}$ is randomly sampled from the Gaussian distribution $\mathcal{N}(0,1)$ and then kept frozen. Thus, we explore the influences of different latent codes by changing their initialization seeds. From Table IX, we discover that although different latent code initialization seeds on the same architecture may result in different initial gradient matching losses, the differences are still very minor (less than $10^{-6}$ for CIFAR-10 and less than $10^{-3}$ for ImageNet), and the impacts on the final PSNR results are limited (less than $0.5$ dB for CIFAR-10 and less than $0.4$ dB for ImageNet). However, when using different architectures, both these two values will change dramatically, and the differences are much greater than those brought by changing the latent code initialization seeds. Hence, GI-NAS is also robust to latent codes initialization.
@@
TABLE IX: Results when attacking the same batch by different networks using different latent code initialization seeds on CIFAR-10 ($32\times 32$) and ImageNet ($256\times 256$) with the default batch size $B=4$. Note that $\sigma_{loss}$ and $\sigma_{psnr}$ are respectively the standard deviations of the initial gradient matching losses and PSNR results on each network.

### v3

@@ IV-B2 Optimal Model Selection
where $\mathcal{L}_{grad}(\cdot)$ is the gradient matching loss, $\mathcal{T}(\cdot)$ is the estimated gradient transformation [11], and $\mathbf{\phi}_{r}$ are the parameters of $G_{r}$. Here we introduce $\mathcal{T}(\cdot)$ to estimate the gradient transformation following the previous defense auditing work [11], since the victims may apply defense strategies to $\mathbf{g}$ (e.g., Gradient Clipping [32]) and only release disrupted forms of gradients. Empirically, the model that can perform the best at the start and stand out from numerous candidates is likely to have better implicit architectural priors with respect to the private batch. This is because the model would have a superior optimization starting point (i.e., a smaller initial loss) due to its inherent structural advantages. We further provide extensive experimental evidence that such a metric highly correlates with the real attack performance in Section V-D. Therefore, we regard the model that achieves the minimal initial loss as our best model $G_{opt}$ and update our selection during the traversal. Since only the initial loss is calculated and no back-propagation is involved, this search process is training-free and hence computationally efficient.
IV-C Stage 2: Private Batch Recovery via the Optimal Model
Similar to Section IV-B, we outline the key components and objectives of this stage as follows:
  • Input for this stage: The searched optimal architecture $G_{opt}$ from Stage 1, the randomly initialized network parameters $\gamma_{1}$, real gradients $\mathbf{g}$, and the fixed latent code $\mathbf{z_{0}}$ from Stage 1.
  • Output for this stage: The final parameters $\mathbf{\gamma^{*}}$, and the reconstructed private data $\mathbf{\hat{x}^{*}}=G_{opt}(\mathbf{z_{0}};\mathbf{\gamma^{*}})$.
  • Purpose for this stage: In this stage, we iteratively optimize the parameters of the selected model $G_{opt}$ to recover the sensitive data.
@@ V-A Experimental Setup
State-of-the-art Baselines for Comparison. We implement the following gradient inversion baselines: (1) IG (Inverting Gradients) [10]: pixel-level reconstruction with angle-based loss function; (2) GI (GradInversion) [17]: realizing batch-level restoration via multiple regularization priors; (3) GGL (Generative Gradient Leakage) [11]: employing strong GAN priors to produce high-fidelity images under severe defense strategies; (4) GIAS (Gradient Inversion in Alternative Space) [18]: searching the optimal latent code while optimizing in the generator parameter space; (5) GIFD (Gradient Inversion over Feature Domain) [12]: leveraging intermediate layer optimization for gradient inversion to further exploit pre-trained GAN priors; (6) GION (Gradient Inversion via Over-parameterized Networks) [13]: designing an over-parameterized convolutional network with excessive parameters and employing a fixed network architecture as implicit regularization. Note that when implementing GAN-based methods such as GGL and GIAS, we adopt BigGAN [70] pre-trained on ImageNet, which may result in mismatched priors when the target data is from CIFAR-10. This is because there is an inherent distribution bias between the ImageNet and CIFAR-10 datasets. However, such mismatched priors can be very common under realistic FL scenarios as mentioned in previous works [13, 12].
Quantitative Metrics. We utilize four metrics to measure the quality of reconstruction images: (1) Peak Signal-to-Noise Ratio (PSNR) $\uparrow$; (2) Structural Similarity Index Measure (SSIM) $\uparrow$; (3) Feature Similarity Index Measure (FSIM) $\uparrow$; (4) Learned Perceptual Image Patch Similarity (LPIPS) $\downarrow$ [71]. Note that $\downarrow$” indicates that the lower the metric, the better the attack performance while $\uparrow$” indicates that the higher the metric, the better the attack performance.
@@ V-B Choices of search size $n$ and network depth $t$
To obtain the best reconstruction performance, we need to carefully decide on the values of search size $n$ and network depth $t$. Note that both these two hyper-parameters reflect the degree of NAS, and there is a trade-off between under-fitting and over-fitting when choosing their values. Thus, we randomly select images (strictly having no intersections with the images used in the remaining experiments of this paper) from ImageNet and test the attack results under different values of $n$ and $t$ with the default batch size $B=4$. From Table I and Table II, we find that when $n$ and $t$ are small (e.g., $n=100$ and $t=3$), increasing any one of them indeed improves the attacks, as more optimal alternatives are provided or the models are more complicated to tackle the under-fitting. However, when $n$ and $t$ are large (e.g., $n=7000$ and $t=6$), further increasing any one of them cannot improve the attacks and will conversely lead to performance degradation due to the over-fitting. This is because candidates that focus too much on the minimal initial loss while neglecting the fundamental patterns of data recovery may be included in the search space and eventually selected. Thus, we adopt $n=5000$ and $t=5$ in the remaining experiments of this paper, as they strike a better balance and acquire the best results in Table I and Table II.
V-C Comparison with State-of-the-art Methods
@@
Dataset Metric IG GI GGL GIAS GIFD GION GI-NAS CIFAR-10 PSNR $\uparrow$ 16.3188 15.4613 12.4938 17.3687 18.2325 30.8652 35.9883 SSIM $\uparrow$ 0.5710 0.5127 0.3256 0.6239 0.6561 0.9918 0.9983 FSIM $\uparrow$ 0.7564 0.7311 0.6029 0.7800 0.8025 0.9960 0.9991 LPIPS $\downarrow$ 0.4410 0.4878 0.5992 0.4056 0.3606 0.0035 0.0009 ImageNet PSNR $\uparrow$ 7.9419 8.5070 11.6255 10.0602 9.8512 21.9942 23.2578 SSIM $\uparrow$ 0.0815 0.1157 0.2586 0.2408 0.2475 0.6188 0.6848 FSIM $\uparrow$ 0.5269 0.5299 0.5924 0.5719 0.5767 0.8198 0.8513 LPIPS $\downarrow$ 0.7194 0.7168 0.6152 0.6563 0.6561 0.4605 0.3952
Firstly, we compare GI-NAS with state-of-the-art gradient inversion methods on CIFAR-10. From Table III, we conclude that GI-NAS achieves the best results with significant performance improvements. For instance, we realize a $5.12$ dB PSNR increase than GION, and our LPIPS value is $74.3\%$ smaller than that of GION. These validate that in contrast to GION that optimizes on a fixed network, our NAS strategy indeed comes into effect and better leverages the implicit architectural priors.
@@
Dataset Metric Batch Size IG GI GGL GIAS GIFD GION GI-NAS Batch Size IG GI GGL GIAS GIFD GION GI-NAS CIFAR-10 PSNR $\uparrow$ 16 10.2647 11.2593 9.9964 11.1952 11.6853 24.8746 30.5315 32 9.5802 10.5863 9.4104 9.8669 11.0898 28.6832 33.0586 SSIM $\uparrow$ 0.2185 0.2431 0.1658 0.2638 0.3222 0.9816 0.9948 0.1595 0.2020 0.1481 0.1724 0.2824 0.9892 0.9974 FSIM $\uparrow$ 0.5254 0.5350 0.4988 0.5569 0.6393 0.9869 0.9965 0.5186 0.5325 0.4995 0.5180 0.6015 0.9882 0.9984 LPIPS $\downarrow$ 0.6254 0.6225 0.6655 0.5804 0.5925 0.0177 0.0035 0.6598 0.6626 0.6922 0.6362 0.6305 0.0218 0.0017 PSNR $\uparrow$ 48 9.5446 10.2000 9.1200 9.6376 10.3709 29.5312 38.7054 96 9.2363 9.9500 8.8575 9.2982 10.3227 28.9723 31.8026 SSIM $\uparrow$ 0.1574 0.1697 0.1552 0.1652 0.2460 0.9932 0.9983 0.1402 0.1643 0.1516 0.2180 0.2389 0.9878 0.9941 FSIM $\uparrow$ 0.5112 0.5167 0.4869 0.5126 0.5824 0.9949 0.9987 0.4961 0.5238 0.4801 0.5365 0.5775 0.9865 0.9941 LPIPS $\downarrow$ 0.6660 0.6571 0.7213 0.6554 0.6504 0.0072 0.0020 0.6865 0.6593 0.7248 0.7297 0.6476 0.0227 0.0071 ImageNet PSNR $\uparrow$ 8 7.6011 8.3419 11.5454 9.6216 8.8377 20.4841 20.8451 16 7.7991 8.0233 11.4766 9.2563 8.8755 20.3078 21.4267 SSIM $\uparrow$ 0.0762 0.1113 0.2571 0.2218 0.2095 0.5681 0.6076 0.0704 0.0952 0.2561 0.2189 0.2193 0.5507 0.6141 FSIM $\uparrow$ 0.5103 0.5365 0.5942 0.5765 0.5726 0.7908 0.8105 0.5077 0.5165 0.5872 0.5832 0.5629 0.7840 0.8157 LPIPS $\downarrow$ 0.7374 0.7251 0.6152 0.6621 0.6701 0.5162 0.4613 0.7424 0.7362 0.6165 0.6656 0.6718 0.5313 0.4613 PSNR $\uparrow$ 24 7.1352 7.9079 10.9910 9.6601 9.2921 20.6845 21.7933 32 7.1085 7.8752 10.8987 10.0563 8.4252 16.1275 20.0011 SSIM $\uparrow$ 0.0549 0.0892 0.2583 0.2270 0.2421 0.5868 0.6236 0.0554 0.0763 0.2558 0.2361 0.2013 0.3415 0.5485 FSIM $\uparrow$ 0.4794 0.4943 0.5811 0.5618 0.5642 0.7978 0.8163 0.4940 0.4951 0.5889 0.5762 0.5545 0.6823 0.7819 LPIPS $\downarrow$ 0.7540 0.7475 0.6261 0.6667 0.6651 0.5219 0.4651 0.7493 0.7427 0.6296 0.6674 0.6853 0.6291 0.5265
@@
Dataset Metric Defense IG GI GGL GIAS GIFD GION GI-NAS Defense IG GI GGL GIAS GIFD GION GI-NAS CIFAR-10 PSNR $\uparrow$ Gaussian Noise 8.8332 8.2797 10.3947 9.8146 9.8441 9.7658 10.5192 Gradient Sparsification 9.6183 10.9486 10.4323 12.1328 14.0772 30.3518 39.8739 SSIM $\uparrow$ 0.1177 0.0976 0.1856 0.1301 0.1734 0.0280 0.0375 0.1592 0.2168 0.1862 0.3227 0.4492 0.9920 0.9987 FSIM $\uparrow$ 0.4844 0.4806 0.5123 0.4843 0.5481 0.5945 0.6300 0.5151 0.5329 0.5223 0.6015 0.6984 0.9951 0.9987 LPIPS $\downarrow$ 0.7258 0.7077 0.6255 0.7129 0.7330 0.6752 0.6740 0.6895 0.6855 0.6298 0.5582 0.5292 0.0047 0.0014 PSNR $\uparrow$ Gradient Clipping 16.6203 11.5888 10.1373 17.9859 19.0230 32.4744 36.2490 Soteria 9.4103 10.2939 10.4217 15.5825 17.6231 31.1003 34.1804 SSIM $\uparrow$ 0.5952 0.2851 0.1760 0.6703 0.7050 0.9933 0.9985 0.1404 0.1707 0.1889 0.5405 0.6319 0.9900 0.9973 FSIM $\uparrow$ 0.7464 0.4976 0.5070 0.7751 0.8237 0.9966 0.9992 0.5086 0.5086 0.5144 0.7133 0.7834 0.9952 0.9986 LPIPS $\downarrow$ 0.3811 0.6413 0.6205 0.3054 0.3313 0.0029 0.0008 0.7008 0.6940 0.6335 0.3748 0.3654 0.0038 0.0014 ImageNet PSNR $\uparrow$ Gaussian Noise 7.7020 7.4553 9.7381 6.8520 8.3526 9.2449 11.2567 Gradient Sparsification 8.6540 8.6519 9.8019 9.7894 9.8620 13.7358 15.4331 SSIM $\uparrow$ 0.0311 0.0197 0.2319 0.1045 0.1618 0.0277 0.0416 0.0669 0.0627 0.2315 0.2515 0.2424 0.2387 0.3154 FSIM $\uparrow$ 0.4518 0.3804 0.5659 0.4562 0.5091 0.5714 0.5846 0.5073 0.4696 0.5624 0.5664 0.5688 0.6365 0.6786 LPIPS $\downarrow$ 0.7775 0.8220 0.6639 0.7282 0.7085 0.8033 0.7805 0.7480 0.7681 0.6589 0.6604 0.6583 0.6909 0.6535 PSNR $\uparrow$ Gradient Clipping 8.0690 7.2342 11.3215 9.7556 9.5759 22.2016 23.2074 Soteria 6.5675 6.6693 11.5690 9.7588 9.0612 22.9617 23.7917 SSIM $\uparrow$ 0.0881 0.0659 0.2531 0.2547 0.2390 0.6266 0.7023 0.0323 0.0382 0.2623 0.2425 0.2166 0.6698 0.7143 FSIM $\uparrow$ 0.5362 0.5434 0.5900 0.5794 0.5731 0.8229 0.8635 0.4929 0.4454 0.5964 0.5697 0.5738 0.8430 0.8658 LPIPS $\downarrow$ 0.7227 0.7230 0.6117 0.6485 0.6653 0.4603 0.3743 0.7506 0.7703 0.6031 0.6631 0.6660 0.4253 0.3655
@@ V-D Further Analysis
Effectiveness of Our Training-Free Search Metric. We provide extensive empirical evidence that our training-free search metric (i.e., the initial gradient matching loss) highly negatively correlates with the real attack performance. Specifically, we attack the same private batch using models with different architectures and report the PSNR performance as well as the training-free search metric of each model both on CIFAR-10 and ImageNet. Following previous training-free NAS methods [61, 62, 63], we calculate the correlation indicators (i.e., Kendall’s $\tau$ [73], Pearson Coefficient [74], Spearman Coefficient [75]) to quantitatively measure the relevance. To be more convincing, we introduce Gaussian Noise $\mathcal{N}\sim(0,1)$ and Uniform Noise $\mathcal{U}\sim(0,1)$ for comparison, where random noises sampled from $\mathcal{N}\sim(0,1)$ and $\mathcal{U}\sim(0,1)$ are regarded as the training-free search metrics. Through this comparative experiment design, we evaluate the performance when using our search method and randomly selecting structures based on Gaussian Noise and Uniform Noise. From Figure 5 and Table VI, we conclude that a smaller initial gradient matching loss is more likely to result in better PSNR performance, while the random noises are far from similar effects. Our method significantly outperforms the two random guessing methods based on Gaussian Noise and Uniform Noise. Thus, utilizing the initial gradient matching loss as the training-free search metric is reasonable and meaningful.
@@
Dataset Metric IG GI GGL GIAS GIFD GION GI-NAS CIFAR-10 NAS Time - - - - - - 6.6 Total Time 46.1 32.9 131.4 214.8 30.2 46.9 57.3 PSNR 16.3188 15.4613 12.4938 17.3687 18.2325 30.8652 35.9883 ImageNet NAS Time - - - - - - 10.9 Total Time 291.1 273.7 244.6 343.2 70.6 58.1 84.4 PSNR 7.9419 8.5070 11.6255 10.0602 9.8512 21.9942 23.2578
TABLE VIII: PSNR results of GION and different variants of GI-NAS on CIFAR-10 ($32\times 32$) and ImageNet ($256\times 256$) with the default batch size $B=4$.
TABLE IX: Attack performance when increasing the training iterations for the search metric on ImageNet ($256\times 256$) with the default batch size $B=4$. Note that the training iteration of 00 indicates the use of our training-free search metric.
Search Metric Training Iteration PSNR $\uparrow$ SSIM $\uparrow$ FSIM $\uparrow$ LPIPS $\downarrow$ 0 (Training-Free Search) 23.2578 0.6848 0.8513 0.3952 5 22.0149 0.6732 0.8545 0.3961 10 20.9854 0.6115 0.8393 0.4511 15 21.0370 0.6260 0.8389 0.4358
TABLE X: PSNR results of GI-NAS and state-of-the-art gradient inversion methods when attacking more FL global models (e.g., LeNet-Zhu [9], ConvNet-32 [10]) on ImageNet ($256\times 256$) with the default batch size $B=4$.
FL Global Model IG GI GGL GIAS GIFD GION GI-NAS LeNet-Zhu [9] 7.7175 6.9530 11.8195 16.5229 17.5620 19.2478 21.1847 ConvNet-32 [10] 10.9081 10.8886 10.4048 8.6696 24.3937 23.9602 25.7824 ResNet-50 [68] 10.9113 10.8914 9.5354 11.0040 9.3521 9.2080 11.2320 ViT-Base [76] 9.7616 8.4317 9.8555 9.9113 10.2516 11.5427 12.6151 ViT-Large [76] 9.6828 8.8500 9.8445 9.9554 10.3850 11.6536 12.4354 ViT-Huge [76] 10.3052 9.1382 10.5788 10.5919 10.4953 11.7839 13.0380
@@
Ablation Study on Search Type. In Table VIII, we report the performance of GION and different variants of GI-NAS. GI-NAS optimizes on a fixed over-parameterized network with the search size $n=1$, which means that it is essentially the same as GION. GI-NAS only searches the skip connection patterns while GI-NAS only searches the upsampling modules. We observe that the performance of GI-NAS is very close to the performance of GION, as both of them adopt a changeless network architecture. GI-NAS and GI-NAS improve the attacks on the basis of GION or GI-NAS, which validates the contribution of each search type. GI-NAS combines the above two search types and thus performs the best among all the variants.
Ablation Study on Search Metric Training Iteration. GI-NAS utilizes the initial gradient matching loss as the training-free search metric, which implies that its training iteration for each candidate is 00. To explore the rationality behind this design, we gradually increase the training iterations for the search metric. From Table IX, we conclude that further increasing the training iterations cannot improve the attacks and will instead increase the time costs. Hence, utilizing the initial gradient matching loss to find the optimal models is reasonable and computationally efficient.
Generalizability to More FL Global Models. We show the performance of different methods when attacking more FL global models (e.g., LeNet-Zhu [9], ConvNet-32 [10]) in Table X. We note that GI-NAS realizes the best reconstruction results in all the tested cases. These results further demonstrate the reliability of GI-NAS. We also observe significant variations in attack results across different FL global models, even when applying the same gradient inversion method (e.g., a $7.85$ dB PSNR drop when switching from LeNet-Zhu [9] to ConvNet-32 [10] on GIAS). This implies that future study may have a deeper look into the correlation between the gradient inversion robustness and the architecture of FL global model, and thus design more securing defense strategies from the perspective of adaptive network architecture choice.
@@
TABLE XI: Results when attacking the same batch by different networks using different parameter initialization seeds on CIFAR-10 ($32\times 32$) and ImageNet ($256\times 256$) with the default batch size $B=4$. The latent code initialization seeds are fixed for all the models. $G_{C1}\sim G_{C3}$ and $G_{I1}\sim G_{I3}$ are randomly sampled from the model search space, while $Seed_{C1}\sim Seed_{C3}$ and $Seed_{I1}\sim Seed_{I3}$ are randomly selected and are different from each other. Note that $\sigma_{loss}$ and $\sigma_{psnr}$ are respectively the standard deviations of the initial gradient matching losses and PSNR results on each network.